The length of a rectangle is 3 cm more than twice its width. If the perimeter of the rectangle is 36 cm, find the dimensions of the rectangle
Answer by ikleyn(53765)


Let the width is  x cm.
Then the length is  2x+3 cm.

The equation is

2x + 2*(2x + 3) = 36.

Simplify and solve it:

2x + 4x + 6 = 36,

6x = 36-6,

6x = 30.

x = 5 cm (the width).
2x+3 = 10+3 = 13 cm (the length).